# Safe Pareto Improvements for Expected Utility Maximizers in Program Games

### v4

@@
Abstract
Agents in mixed-motive coordination problems such as Chicken may fail to coordinate on a Pareto-efficient outcome. Safe Pareto improvements (SPIs) were originally proposed to mitigate miscoordination in cases where players lack probabilistic beliefs as to how their agents will play a game; agents are instructed to behave so as to guarantee a Pareto improvement on how they would play by default. More generally, SPIs may be defined as transformations of strategy profiles such that all players are necessarily better off under the transformed profile. In this work, we investigate the extent to which SPIs can reduce downsides of miscoordination between expected utility-maximizing agents. We consider games in which players submit computer programs that can condition their decisions on each other’s code, and use this property to construct SPIs using programs capable of renegotiation. We first show that under mild conditions on players’ beliefs, each player always prefers to use renegotiation. Next, we show that under similar assumptions, each player always prefers to be willing to renegotiate at least to the point at which they receive the lowest payoff they can attain in any efficient outcome. Thus subjectively optimal play guarantees players at least these payoffs, without the need for coordination on specific Pareto improvements. Lastly, we prove that renegotiation does not guarantee players any improvements on this bound.
1 Introduction
@@
We will explore how the ability of AI systems to condition their decisions on each other’s inner workings could ameliorate miscoordination in bargaining problems. The literature on program equilibrium has shown how games played by computer programs that can read each other’s source code admit more cooperative equilibria in other challenges for cooperation such as the Prisoner’s Dilemma (Tennenholtz, 2004; LaVictoire et al., 2014; Oesterheld, 2019). Safe Pareto improvements (SPIs) (Oesterheld and Conitzer, 2022) were proposed as a mitigation for inefficiencies in settings where players have delegates play a game on their behalf, and have Knightian uncertainty (i.e., lack probabilistic beliefs; (Knight, 1921)) about how their delegates will play. Under an SPI, players instruct their delegates to change their default policies so as to guarantee Pareto improvement on the default outcome. For example, consider two parties $A$ and $B$ who would by default go to war over some territory. They might instruct their delegates to, instead, accept the outcome of a lottery that allocates the territory to $A$ with the probability that $A$ would have won the war.
We will consider the extent to which SPIs can mitigate inefficiencies from miscoordination when (i) players do have probabilistic beliefs and maximize subjective expected utility and (ii) games are played by computer programs that can condition on their counterparts’ source code. Our contributions are as follows:
  1. 1.
@@
    We construct SPIs in the program game setting using programs that renegotiate. Such programs have a “default” program; check if their default played against their counterparts’ defaults results in an inefficient outcome; and, if so, call a renegotiation routine in an attempt to Pareto-improve on the default outcome. We examine when renegotiation would be used by players who optimize expected utility given their beliefs about what their counterparts will do (i.e., in subjective equilibrium (Kalai and Lehrer, 1995)). Under mild assumptions on players’ beliefs, we show that SPIs are always used in subjective equilibrium (Propositions 1 and 2).
  2. 2.
@@
    We show that due to the ability to renegotiate, under mild assumptions on players’ beliefs, players always weakly prefer programs that guarantee they receive at least the lowest payoff they can obtain on the Pareto frontier (Theorem 3). Following Rabin (1991), we refer to this payoff profile as the Pareto meet minimum (PMM). Thus we provide for this setting a (partial) solution to the “SPI selection problem” identified by Oesterheld and Conitzer (2022) (hereafter, “OC”), i.e., the problem that players must coordinate among SPIs in order to Pareto-improve on default outcomes. Moreover, our bound is tight: In mixed-motive games, it is always possible to find subjective equilibria in which players fail to Pareto-improve on the PMM, even using iterated renegotiation (Proposition 5). The intuition for this is simple: The PMM is the most efficient point from which all players still have a chance to renegotiate to their ideal outcome. And given some beliefs, it will be subjectively optimal not to agree to anything less than the ideal outcome, despite the risk of failing to agree on a Pareto-improvement.
2 Related Work
@@
We build on program games, where computer players condition their actions on each other’s source code. Prior work has shown that the ability of computer-based agents to condition their decisions on their counterparts’ programs can enable more efficient equilibria (McAfee, 1984; Howard, 1988; Rubinstein, 1998; Tennenholtz, 2004; LaVictoire et al., 2014; Critch, 2019; Oesterheld, 2019; DiGiovanni and Clifton, 2023). For example, McAfee (1984)’s program “If other player’s code == my code: Cooperate; Else: Defect” is a Nash equilibrium of the program game version of the one-shot Prisoner’s Dilemma in which both players cooperate. (See also the literature on commitment games, e.g., Kalai et al. (2010); Forges (2013).) However, this literature focuses on the Nash equilibria of program games, rather than studying failure to coordinate on a Nash equilibrium as we do here.
@@
There are large theoretical and empirical literatures on how agents might coordinate in complete information bargaining problems (see Schuessler and Van der Rijt (2019) and references therein). Most closely related to this paper is the literature on whether communication before playing a simultaneous-move game of complete information can improve coordination (Farrell, 1987; Rabin, 1991; Farrell and Rabin, 1996; Crawford, 2017; He et al., 2019). Rabin (1991) considers solution concepts for games with pre-play communication called negotiated equilibrium (NGE) and negotiated rationalizability (NGR), where NGE assumes that players know their counterpart’s strategies exactly (up to randomization). Rabin shows that under NGE players are guaranteed at least their PMM payoff in bargaining problems, whereas under NGR they are not. NGR is closer to the notion of subjective equilibrium used in our paper, which allows players to have possibly-inaccurate beliefs about what programs their counterparts will use. Santos (2000) shows results analogous to Rabin (1991)’s under cheap talk with alternating (rather than simultaneous) announcements. Following a suggestion by Yudkowsky (2013), Diffractor (2022) develops a cooperative bargaining solution that uses the PMM as the disagreement point. Finally, OC proposed safe Pareto improvements for mitigating inefficiencies from coordination failures. We discuss OC and its connections to the present work at greater length in Section 3.2.
3 Miscoordination and Safe Pareto Improvements in Program Games
In this section, we introduce the program games framework and subjective equilibrium, the solution concept that is our focus in this paper. Then we review OC’s safe Pareto improvements, and show how they can be constructed in our setting using renegotiation. Appendix A contains a table summarizing the notation used in this section and Section 4.
3.1 Setup: Program games and subjective equilibrium
@@
A program game $G(\mathcal{P})$ is a game in which a strategy is a program that maps the profile of other players’ programs to an action in $G$ [^1] We restrict to deterministic programs for ease of exposition; the extension to probabilistic programs, as in, e.g., Kalai et al. (2010), is straightforward. This way, each player’s program implements a commitment to an action conditional on the others’ programs. Here, $\mathcal{P}=\bigtimes_{i=1}^{{n}}P_{i}$, where $P_{i}$ is a set of computable functions from ×$\bigtimes_{j\neq i}P_{j}$ to $A_{i}$. We assume that all programs in $P_{i}$ halt against all program profiles in ×$\bigtimes_{j\neq i}P_{j}$, for all $i$, as is standard in program game literature (see, e.g., Tennenholtz (2004); Oesterheld (2019); Oesterheld and Conitzer (2022)). (Each $P_{i}$ can be viewed as player $i$’s “default” program set, which we will extend in Section 3.3 with a set of programs that have a special structure.)
Player $i$’s program is $p_{i}\in P_{i}$. For program profiles $\textbf{p}=(p_{i})_{i=1}^{{n}}$ and $\textbf{p}_{-i}=(p_{j})_{j\neq i}$, abusing notation, let the action profile played in the base game by players with a given program profile be $\boldsymbol{a}(\textbf{p})=(p_{i}(\textbf{p}_{-i}))_{i=1}^{{n}}$. After all programs are simultaneously submitted, the induced action profile $\boldsymbol{a}(\textbf{p})$ is played in $G$. Thus the payoff for player $i$ in $G(\mathcal{P})$ resulting from the program profile p is $U_{i}(\textbf{p})=u_{i}(\boldsymbol{a}(\textbf{p}))$.
To capture the possibility of miscoordination, we do not assume a Nash equilibrium is played. Instead, each player $i$ has beliefs as to what program profile $\textbf{p}_{-i}$ the other players will use, distributed according to a probability distribution $\beta_{i}$ (whose support may be a superset of ×$\bigtimes_{j\neq i}P_{j}$). [^2] Allowing for $\beta_{i}$ to be supported on a superset of ×$\bigtimes_{j\neq i}P_{j}$ will be important when we consider extensions of players’ program sets with SPIs in Section 3.3. Then, a subjective equilibrium (Kalai and Lehrer, 1995) is a profile of programs and beliefs such that each player’s program maximizes expected utility with respect to their beliefs:
@@
Let $\textbf{p}^{*}=(p_{i}^{*})_{i=1}^{{n}}$ and $\boldsymbol{\beta}=(\beta_{i})_{i=1}^{{n}}$ be profiles of programs and beliefs, respectively, in $G(\mathcal{P})$. We say $(\textbf{p}^{*},\boldsymbol{\beta})$ is a subjective equilibrium of $G(\mathcal{P})$ if, for all $i$,
$_{i}^{*}\in\operatorname*{arg\,max}_{p_{i}\in P_{i}}\mathbb{E}_{\textbf{p}_{-i}\sim\beta_{i}}U_{i}(\textbf{p}).$ p ) .
Subjective equilibrium is, of course, a weaker solution concept than Nash equilibrium (or even rationalizable strategies (Bernheim, 1984; Pearce, 1984)). The results in this paper that follow will be stronger than showing that a given strategy is used in some subjective equilibrium. Instead, we will construct strategies such that, for any beliefs players might have under some assumptions, and any program profile they consider using, our strategies are individually preferred by players over that program profile — and are thus used in a subjective equilibrium associated with those beliefs. Therefore, considering subjective equilibrium will make our results stronger than if we had assumed players’ beliefs satisfied a Nash equilibrium assumption.
@@
(Miscoordination in subjective equilibrium) Suppose two principals delegate to AI assistants to negotiate on their behalf over the time for a meeting. Call this the Scheduling Game (Table 1). The principals meet if and only if the AIs agree on one of three possible time slots. Each principal $i$ most prefers slot $i$, but would rather meet at slot 3 than not at all. Suppose each player $i$ thinks $j$ is sufficiently likely to use [^3] We will abuse notation throughout by writing “$p_{i}=\langle\text{pseudocode for }p_{i}\rangle$” to describe programs $p_{i}$. $p_{j}^{C}=$ “Slot $i$ if other player’s code == ‘always Slot $i$’; Else: Slot $j$”, which can only be exploited by $p_{i}^{D}=$ “always Slot $i$”. Each player might believe the other is likely to use $p_{j}^{C}$ because it can both exploit programs that yield to its demand and avoid miscoordinating with $p_{i}^{D}$. Then it is subjectively optimal for each to submit $p_{i}^{D}$. The pair of programs $(p_{1}^{D},p_{2}^{D})$ played in a subjective equilibrium under these beliefs results in the maximally inefficient (Slot 1, Slot 2) outcome.
Table 1: Payoff matrix for the Scheduling Game.
@@
3.2 Review of safe Pareto improvements
@@
[Introduced by OC as “strict SPI”] A subset game $G^{\textnormal{s}}$ of $G$ is a strict payoff-transformation SPI on $G$ if, for any distribution over $\pi\in\Pi$, we have that $P[\textbf{u}(\pi(G^{\textnormal{s}}))\succeq\textbf{u}(\pi(G))]=1$ and $P[u_{i^{\prime}}(\pi(G^{\textnormal{s}}))>u_{i^{\prime}}(\pi(G))\text{ for some $i^{\prime}$}]>0$.
Note that OC refer to the object defined in Definition 2 as just “strict SPI.” However, they note that probability-1 Pareto improvements on representatives’ default strategies can be achieved with other kinds of instructions besides having representatives play a game with transformed payoffs. [^4] OC, pg. 14 (but using our notation): “In principle, Theorem 3 [which shows that payoff-transformation SPIs require a certain kind of “outcome correspondence” between $G$ and $G^{\textnormal{s}}$] does not hinge on $\pi(G)$ and $\pi(G^{\textnormal{s}})$ resulting from playing games. An analogous result holds for any random variables over $A_{1}\times A_{2}$ and $A_{1}^{\textnormal{s}}\times A_{2}^{\textnormal{s}}$.” Thus in this paper we define SPIs to be general transformations of strategy profiles that guarantee Pareto improvement:
Definition 3.
For a program game $G(\mathcal{P})$, let $\textbf{f}:\mathcal{P}\to\mathcal{P}^{\prime}$ be a function of program profiles, written $\textbf{f}(\textbf{p})=(f_{i}(p_{i}))_{i=1}^{{n}}$, for some joint program space $\mathcal{P}^{\prime}=\bigtimes_{i=1}^{{n}}P_{i}^{\prime}$. [^5] The assumption above that programs halt against each other extends to $\mathcal{P}^{\prime}$. Then f is an SPI for $G(\mathcal{P})$ if, for all program profiles p, we have $\textbf{U}(\textbf{f}(\textbf{p}))\succeq\textbf{U}(\textbf{p})$; and for some program profile p, there is some $i^{\prime}$ such that $U_{i^{\prime}}(\textbf{f}(\textbf{p}))>U_{i^{\prime}}(\textbf{p})$.
3.3 Constructing SPIs via renegotiation
A natural approach to constructing an SPI is to construct programs that, when they are all used against each other, map the action profile returned by default programs to a Pareto improvement whenever the default programs would have otherwise miscoordinated (i.e., the action profile is inefficient). We call this “renegotiation.” [^6] Compare to section “Safe Pareto improvements under improved coordination” in OC.
@@ 3.3 Constructing SPIs via renegotiation
We jointly define the spaces of renegotiation programs $P_{i}^{\textnormal{rn}}(\mathbf{rn}^{i})$ for $i=1,\dots,{n}$ as those programs with the structure of Algorithm 1, for some:
@@
(Note that the definition of Algorithm 1 for a given player $i$ references the sets of programs given by Algorithm 1 for the other players $j\neq i$, so this definition is not circular.) For any program profile $\textbf{p}\in\bigtimes_{i=1}^{{n}}P_{i}^{\textnormal{rn}}(\mathbf{rn}^{i})$ and any renegotiation function $\mathbf{rn}$, we write ${\textbf{p}}^{\textnormal{def}}=(p_{i}^{\textnormal{def}})_{i=1}^{{n}}$ and $\mathbf{rn}(\boldsymbol{a})=(\textnormal{rn}_{i}(\boldsymbol{a}))_{i=1}^{{n}}$.
Renegotiation programs work as follows: Consider the “default outcome,” the action profile given by all players’ default programs (if players all use renegotiation programs). Against any program profile $\textbf{p}_{-i}$ such that the players’ renegotiation functions (if any) don’t all return the same Pareto improvement on the default outcome, $p_{i}\in P_{i}^{\textnormal{rn}}(\mathbf{rn}^{i})$ plays according to its default program $p_{i}^{\textnormal{def}}$. Against a program profile that is willing to renegotiate to the same Pareto improvement, however, $p_{i}$ plays its part of the Pareto-improved outcome.
@@
1:Counterpart program profile $\textbf{p}_{-i}$
2:if $\textbf{p}_{-i}\in\bigtimes_{j\neq i}P_{j}^{\textnormal{rn}}(\mathbf{rn}^{j})$ for some $(\mathbf{rn}^{j})_{j\neq i}\in\bigtimes_{j\neq i}\mathcal{R}$ then $\triangleright$ Check that $\textbf{p}_{-i}$ renegotiate
@@
4:     if $\mathbf{rn}^{i}(\widehat{\boldsymbol{a}})=\mathbf{rn}^{j}(\widehat{\boldsymbol{a}})$ for all $j\neq i$ then
@@
9:     return $p_{i}^{\textnormal{def}}(\textbf{p}_{-i})$triangleright Play default
@@
Let $\mathbf{rn}$ be a renegotiation function. For $i=1,\dots,{n}$, define $f_{i}:P_{i}\rightarrow P_{i}^{\textnormal{rn}}(\mathbf{rn})$ such that, for each $p_{i}\in P_{i}$, $f_{i}(p_{i})$ is of the form given in Algorithm 1 with ${f_{i}(p_{i})}^{\textnormal{def}}=p_{i}$. Then, the function $\textbf{f}:\textbf{p}\mapsto(f_{i}(p_{i}))_{i=1}^{{n}}$ is an SPI.
@@
3.4 Incentives to renegotiate
When is the use of renegotiation guaranteed in subjective equilibrium in program games? SPIs by definition make all players (weakly) better off ex post, but it remains to show that players prefer to renegotiate ex ante. It is plausible that, all else equal, players prefer strategies that admit more opportunities for coordination. So, suppose players always prefer a renegotiation program over a non-renegotiation program if their expected utility is unchanged. Then Proposition 2 shows that, under a mild assumption on players’ beliefs, each player always prefers to transform their default program into some renegotiation program. That is, each player $i$ always prefers to use a program in $P_{i}^{\textnormal{rn}}=\bigcup_{\mathbf{rn}}P_{i}^{\textnormal{rn}}(\mathbf{rn})$.
For this result, we assume (Assumption 5) the following holds for any program profile p given by (a) a program used by some player $i$ in subjective equilibrium and (b) programs in the support of player $i$’s beliefs: Suppose that the programs in p don’t renegotiate with each other. Then, a program should respond equivalently to any renegotiation programs as it would to respond to those programs’ defaults. It seems implausible that players would respond differently to renegotiation programs that do not respond differently to them (in particular, “punish” renegotiation), all else equal. (All full proofs are in the appendix.)

### v5

@@
Abstract.
Agents in mixed-motive coordination problems such as Chicken may fail to coordinate on a Pareto-efficient outcome. Safe Pareto improvements (SPIs) were originally proposed to mitigate miscoordination in cases where players lack probabilistic beliefs as to how their agents will play a game; agents are instructed to behave so as to guarantee a Pareto improvement on how they would play by default. More generally, SPIs may be defined as transformations of strategy profiles such that all players are necessarily better off under the transformed profile. In this work, we investigate the extent to which SPIs can reduce downsides of miscoordination between expected utility-maximizing agents. We consider games in which players submit computer programs that can condition their decisions on each other’s code, and use this property to construct SPIs using programs capable of renegotiation. We first show that under mild conditions on players’ beliefs, each player always prefers to use renegotiation. Next, we show that under similar assumptions, each player always prefers to be willing to renegotiate at least to the point at which they receive the lowest payoff they can attain in any efficient outcome. Thus subjectively optimal play guarantees players at least these payoffs, without the need for coordination on specific Pareto improvements.
Key words and phrases:
program equilibrium; bargaining; Pareto efficiency; Cooperative AI
1. Introduction
@@
We will explore how the ability of AI systems to condition their decisions on each other’s inner workings could reduce downsides of miscoordination in bargaining problems. The literature on program equilibrium has shown how games played by computer programs that can read each other’s source code admit more cooperative equilibria in other challenges for cooperation such as the Prisoner’s Dilemma (Tennenholtz, 2004; LaVictoire et al., 2014; Oesterheld, 2019). Safe Pareto improvements (SPIs) (Oesterheld and Conitzer, 2022) were proposed as a mitigation for inefficiencies in settings where players have delegates play a game on their behalf, and have Knightian uncertainty (i.e., lack probabilistic beliefs (Knight, 1921)) about how their delegates will play. Under an SPI, players change their default policies so as to guarantee Pareto improvement on the default outcome. For example, consider two parties $A$ and $B$ who would by default go to war over some territory. They might instruct their delegates to, instead, accept the outcome of a lottery that allocates the territory to $A$ with the probability that $A$ would have won the war.
We will consider the extent to which SPIs can mitigate inefficiencies from miscoordination when (i) players do have probabilistic beliefs and maximize subjective expected utility and (ii) games are played by computer programs that can condition on their counterparts’ source code. Our goal is to establish guarantees against miscoordination in the well-studied program game setting. Relaxations of standard assumptions in this setting — e.g., players can precisely read each other’s programs’ source code, can syntactically verify if a program follows some template (Tennenholtz, 2004), and participate in the program game in the first place — are left to future work. While this is an idealized framework, insights from studying program games could be applied to more realistic interactions between actors with some degree of conditional commitment ability. For example, countries engaging in climate negotiations might write bills that specify when the country would be bound to some policies conditional on the terms of other countries’ bills (Heitzig et al., 2023). And, smart contracts implemented on a blockchain could execute commitments to transactions conditional on other actors’ contracts (Varian, 2010; Sun et al., 2023).
Our contributions are as follows:
  1. (1)
@@
    We construct SPIs in the program game setting using programs that renegotiate. Such programs have a “default” program; check if their default played against their counterparts’ defaults results in an inefficient outcome; and, if so, call a renegotiation routine in an attempt to Pareto-improve on the default outcome. We examine when renegotiation would be used by players who optimize expected utility given their beliefs about what programs their counterparts will use (i.e., in subjective equilibrium (Kalai and Lehrer, 1995)). Under mild assumptions on players’ beliefs, we show that SPIs are always used in subjective equilibrium (Propositions 4 and 5).
  2. (2)
@@
    We show that due to the ability to renegotiate, under mild assumptions on players’ beliefs, players always weakly prefer programs that guarantee them at least the lowest payoff they can obtain on the Pareto frontier (Theorem 9). Following Rabin (1991), we call this payoff profile the Pareto meet minimum (PMM). Thus we provide for this setting a (partial) solution to the “SPI selection problem” identified by Oesterheld and Conitzer (2022) (hereafter, “OC”), i.e., the problem that players must coordinate among SPIs in order to Pareto-improve on default outcomes. The intuition for this is: The PMM is the most efficient point such that, no matter how aggressively the players bargain, no one expects to risk getting a worse deal by being willing to renegotiate to that point. We also show in the appendix that the PMM bound is tight: In mixed-motive games, it is always possible to find subjective equilibria in which players fail to Pareto-improve on the PMM, even using iterated renegotiation (Proposition 5).
2. Related Work
@@
We build on program games, where computer players condition their actions on each other’s source code. Prior work has shown that the ability of computer-based agents to condition their decisions on their counterparts’ programs can enable more efficient equilibria (McAfee, 1984; Howard, 1988; Rubinstein, 1998; Tennenholtz, 2004; LaVictoire et al., 2014; Critch, 2019; Oesterheld, 2019; DiGiovanni and Clifton, 2023). For example, McAfee (1984)’s program “If other player’s code == my code: Cooperate; Else: Defect” is a Nash equilibrium of the program game version of the one-shot Prisoner’s Dilemma in which both players cooperate. (See also the literature on commitment games, e.g., Kalai et al. (2010); Forges (2013).) However, this literature focuses on the Nash equilibria of program games, rather than studying failure to coordinate on a Nash equilibrium as we do.
@@
There are large theoretical and empirical literatures on how agents might coordinate in complete information bargaining problems (see Schuessler and Van der Rijt (2019) and references therein). Most closely related to this paper is the literature on whether communication before playing a simultaneous-move game can improve coordination (Farrell, 1987; Rabin, 1991; Farrell and Rabin, 1996; Crawford, 2017; He et al., 2019). Rabin (1991) considers solution concepts for games with pre-play communication called negotiated equilibrium (NGE) and negotiated rationalizability (NGR), where NGE assumes that players know their counterpart’s strategies exactly (up to randomization). Rabin shows that under NGE players are guaranteed at least their PMM payoff in bargaining problems, whereas under NGR they are not. NGR is closer to the notion of subjective equilibrium used in our paper, which allows players to have possibly-inaccurate beliefs about what programs their counterparts will use. Santos (2000) shows results analogous to Rabin (1991)’s under cheap talk with alternating (rather than simultaneous) announcements. Finally, OC proposed safe Pareto improvements for mitigating inefficiencies from coordination failures. We discuss OC and its connections to the present work at greater length in Section 3.2.
3. Miscoordination and Safe Pareto Improvements in Program Games
In this section, we introduce the program games framework and subjective equilibrium, the solution concept that is our focus in this paper. Then we review OC’s safe Pareto improvements, and show how they can be constructed in our setting using renegotiation. Section 5 contains a table summarizing the notation used in this section and Section 4. Throughout the paper, our formalism will be for games with two players, for ease of exposition. See appendix for full proofs of our results in the more general ${n}$-player formalism. The extension to ${n}$ players doesn’t introduce qualitatively new challenges. Intuitively, since players submit programs independently of each other, we can apply the same arguments to the profile of counterparts for a given player, as we did to the single counterpart in the two-player case.
3.1. Setup: Program Games and Subjective Equilibrium
@@
A program game $G(\mathcal{P})$ is a game in which a strategy is a program that maps the profile of other players’ programs to an action in $G$ [^1] We restrict to deterministic programs for ease of exposition; the extension to probabilistic programs, as in, e.g., Kalai et al. (2010), is straightforward. This way, each player’s program implements a commitment to an action conditional on the others’ programs. Assume the action sets of $G$ are continuous; this is practically without loss of generality, because our program game setting can be extended to a setting where players can use correlated randomization (see, e.g., Kalai et al. (2010)). Here, $\mathcal{P}=P_{1}\times P_{2}$, where $P_{i}$ is a set of computable functions from $P_{j}$ to $A_{i}$. We assume that all programs in $P_{i}$ halt against all programs in $P_{j}$, for each $i$, as is standard in program game literature (see, e.g., Tennenholtz (2004); Oesterheld (2019); Oesterheld and Conitzer (2022)). (Each $P_{i}$ can be viewed as player $i$’s “default” program set, which we will extend in Section 3.2 with a set of programs that have a special structure.)
Player $i$’s program is $p_{i}\in P_{i}$. For a program profile $\textbf{p}=(p_{1},p_{2})$, abusing notation, let the action profile played in the base game by players with a given program profile be $\boldsymbol{a}(\textbf{p})=(p_{1}(p_{2}),p_{2}(p_{1}))$. After all programs are simultaneously submitted, the induced action profile $\boldsymbol{a}(\textbf{p})$ is played in $G$. Thus the payoff for player $i$ in $G(\mathcal{P})$ resulting from the program profile p is $U_{i}(\textbf{p})=u_{i}(\boldsymbol{a}(\textbf{p}))$.
To capture the possibility of miscoordination, we do not assume a Nash equilibrium is played. Instead, each player $i$ has beliefs as to what program $p_{j}$ the other player will use, distributed according to a probability distribution $\beta_{i}$ (whose support may be a superset of $P_{j}$). [^2] Allowing for $\beta_{i}$ to be supported on a superset of $P_{j}$ will be important when we consider extensions of players’ program sets with SPIs in Section 3.2. Then, a subjective equilibrium (Kalai and Lehrer, 1995) is a profile of programs and beliefs such that each player’s program maximizes expected utility with respect to their beliefs:
@@
Let $\textbf{p}^{*}=(p_{1}^{*},p_{2}^{*})$ and $\boldsymbol{\beta}=(\beta_{1},\beta_{2})$ be profiles of programs and beliefs, respectively, in $G(\mathcal{P})$. We say $(\textbf{p}^{*},\boldsymbol{\beta})$ is a subjective equilibrium of $G(\mathcal{P})$ if, for each $i$,
$_{i}^{*}\in\operatorname*{arg\,max}_{p_{i}\in P_{i}}\mathbb{E}_{p_{j}\sim\beta_{i}}U_{i}(\textbf{p}).$ p ) .
Subjective equilibrium is, of course, a weaker solution concept than Nash equilibrium (or even rationalizable strategies (Bernheim, 1984; Pearce, 1984)). The results in this paper that follow will be stronger than showing that a given strategy is used in some subjective equilibrium. Instead, we will construct strategies such that, for any beliefs players might have under some assumptions, and any program profile they consider using, our strategies are individually (weakly) preferred by players over that program profile — and are thus used in a subjective equilibrium associated with those beliefs. Therefore, considering subjective equilibrium will make our results stronger than if we had assumed players’ beliefs satisfied a Nash equilibrium assumption.
@@
(Miscoordination in subjective equilibrium) Suppose two principals delegate to AI assistants to negotiate on their behalf over the time for a meeting. Call this the Scheduling Game (Table 1). The principals meet if and only if the AIs agree on one of three possible time slots. Each principal $i$ most prefers slot $i$, but would rather meet at slot 3 than not at all. Suppose each player $i$ thinks $j$ is sufficiently likely to use [^3] Abusing notation, we write “$p_{i}=\langle\text{pseudocode for }p_{i}\rangle$” to describe programs $p_{i}$. $p_{j}^{C}=$ “Slot $i$ if other player’s code == ‘always Slot $i$’; Else: Slot $j$. Intuitively, this program “demands” the player’s best outcome, except against $p_{i}^{D}=$ “always Slot $i$, which exploits this program. Each player might believe the other is likely to use $p_{j}^{C}$ because it can both exploit programs that yield to its demand and avoid miscoordinating with $p_{i}^{D}$. Then it is subjectively optimal for each player to submit $p_{i}^{D}$. The pair of programs $(p_{1}^{D},p_{2}^{D})$ played in a subjective equilibrium under these beliefs results in the maximally inefficient (Slot 1, Slot 2) outcome.
Table 1. Payoff matrix for the Scheduling Game
@@
3.2. Constructing Safe Pareto Improvements via Renegotiation
@@
For a program game $G(\mathcal{P})$, let $\textbf{f}:\mathcal{P}\to\mathcal{P}^{\prime}$ be a function of program profiles, written $\textbf{f}(\textbf{p})=(f_{1}(p_{1}),f_{2}(p_{2}))$, for some joint program space $\mathcal{P}^{\prime}=P_{1}^{\prime}\times P_{2}^{\prime}$. [^4] The assumption above that programs halt against each other extends to $\mathcal{P}^{\prime}$. Then f is an SPI for $G(\mathcal{P})$ if, for all program profiles p, we have $\textbf{U}(\textbf{f}(\textbf{p}))\succeq\textbf{U}(\textbf{p})$; and for some program profile p, there is some $i^{\prime}$ such that $U_{i^{\prime}}(\textbf{f}(\textbf{p}))>U_{i^{\prime}}(\textbf{p})$.
A natural approach to constructing an SPI is to construct programs that, when they are all used against each other, map the action profile returned by default programs to a Pareto improvement whenever the default programs would have otherwise miscoordinated (i.e., the action profile is inefficient). We call this construction “renegotiation,” and call mappings of action profiles to Pareto improvements “renegotiation functions.” [^5] Compare to section “Safe Pareto improvements under improved coordination” in OC.
@@
We jointly define the spaces of renegotiation programs $P_{i}^{\textnormal{rn}}(\mathbf{rn}^{i})$ for $i=1,2$ as those programs with the structure of Algorithm 1, for some:
@@
(Note that the definition of Algorithm 1 for a given player $i$ references the sets of programs given by Algorithm 1 for the other player $j$, so this definition is not circular.) For any program profile $\textbf{p}\in P_{1}^{\textnormal{rn}}(\mathbf{rn}^{1})\times P_{2}^{\textnormal{rn}}(\mathbf{rn}^{2})$ and any renegotiation function $\mathbf{rn}$, we write ${\textbf{p}}^{\textnormal{def}}=(p_{1}^{\textnormal{def}},p_{2}^{\textnormal{def}})$ and $\mathbf{rn}(\boldsymbol{a})=(\textnormal{rn}_{1}(\boldsymbol{a}),\textnormal{rn}_{2}(\boldsymbol{a}))$.
Renegotiation programs work as follows: Consider the “default outcome,” the action profile given by all players’ default programs if they all use renegotiation programs (line 3). Against any program $p_{j}$ such that the players’ renegotiation functions (if any) don’t all return the same Pareto improvement on the default outcome, $p_{i}\in P_{i}^{\textnormal{rn}}(\mathbf{rn}^{i})$ plays according to its default program $p_{i}^{\textnormal{def}}$ (lines 7 and 9 in Algorithm 1). Against a program profile that is willing to renegotiate to the same Pareto improvement, however, $p_{i}$ plays its part of the Pareto-improved outcome (line 5).
@@
1:Counterpart program $p_{j}$
2:if $p_{j}\in P_{j}^{\textnormal{rn}}(\mathbf{rn}^{j})$ for some $\mathbf{rn}^{j}$ $\in\mathcal{R}$ then $\triangleright$ Check that $p_{j}$ renegotiates
@@
4:     if $\mathbf{rn}^{i}(\widehat{\boldsymbol{a}})=\mathbf{rn}^{j}(\widehat{\boldsymbol{a}})$ then
@@
9:     return $p_{i}^{\textnormal{def}}(p_{j})$triangleright Play default
@@
Let $\mathbf{rn}$ be a renegotiation function. For $i=1,2$, define $f_{i}:P_{i}\rightarrow P_{i}^{\textnormal{rn}}(\mathbf{rn})$ such that, for each $p_{i}\in P_{i}$, $f_{i}(p_{i})$ is of the form given in Algorithm 1 with ${f_{i}(p_{i})}^{\textnormal{def}}=p_{i}$. Then, the function $\textbf{f}:\textbf{p}\mapsto(f_{1}(p_{1}),f_{2}(p_{2}))$ is an SPI.
@@
3.3. Incentives to Renegotiate
When is the use of renegotiation guaranteed in subjective equilibrium in program games? SPIs by definition make all players (weakly) better off ex post, but it remains to show that players prefer to renegotiate ex ante. Intuitively, one might worry that players will choose not to accept a Pareto improvement in order to avoid losing bargaining power.
It is plausible that, all else equal, players prefer strategies that admit more opportunities for coordination. So, suppose players always prefer a renegotiation program over a non-renegotiation program if their expected utility is unchanged. Then Proposition 5 shows that, under a mild assumption on players’ beliefs, each player always prefers to transform their default program into some renegotiation program. That is, each player $i$ always prefers to use a program in $P_{i}^{\textnormal{rn}}=\bigcup_{\mathbf{rn}}P_{i}^{\textnormal{rn}}(\mathbf{rn})$ (so their program profile is in $\mathcal{P}^{\text{rn}}=P_{1}^{\textnormal{rn}}\times P_{2}^{\textnormal{rn}}$).
For this result, we assume (Assumption 1) the following holds for any program profile p given by (a) a program used by some player $i$ in subjective equilibrium and (b) a program in the support of player $i$’s beliefs: If the programs in p don’t renegotiate with each other, then, a program should respond equivalently to any renegotiation program as it would respond to that program’s default. This is because it seems implausible that players would respond differently to renegotiation programs that do not respond differently to them (in particular, “punish” renegotiation), all else equal. (All full proofs are in the appendix.)